# Adaptive Monitoring of Stochastic Fire Front Processes via Information-seeking Predictive Control

### v1

@@ III-B Information-seeking Predictive Control
For each prediction time step $\tau\in\{1,\ldots,T\}$ in the horizon, the agent predicts the fire front state $X_{\tau}$ forward in time using the Bayesian prediction step, based on the transition density $p(X_{\tau}|X_{\tau-1})$ and the (pseudo) posterior belief from the previous time step, $\mathcal{B}_{t+\tau-1|t}(X_{\tau-1}|Z_{1:t+\tau-1|t})$, as shown in Eqs. (6b)-(6c). The constraints in Eqs. (6d)-(6e) arise from the agent dynamics, which are assumed to be deterministic in this work. At the predicted time step $\tau$, given the agent’s predicted state $y_{t+\tau|t}$ and sensing range $O_{t+\tau|t}$, the agent receives the predicted measurement set $Z_{t+\tau|t}$. This set is then used to compute the posterior belief $\mathcal{B}_{t+\tau|t}(X_{\tau}|Z_{1:t+\tau|t})$ via the Bayesian correction step, as shown in Eq. (6f), using the joint likelihood function $p(Z_{t+\tau|t}|X_{\tau},y_{t+\tau|t})$ and the predicted density. This posterior distribution subsequently becomes the prior for the next prediction step, continuing the recursive process. The predicted measurements $Z_{t+\tau|t}$ represent hypothetical observations based on the planned control inputs and the anticipated fire front state. Since actual measurements are only available after executing the control actions, the objective in Eq. (6a) requires taking an expectation over all possible future measurement sequences. This enables informed decision-making by accounting for potential outcomes without executing the corresponding trajectories.
@@
Equations (6b)-(6f) admit no closed form: the model is nonlinear and non-Gaussian with set-valued (multi-object, multi-measurement) observations, so Kalman-type filters are inapplicable. Therefore, the recursion is implemented using Sequential Importance Resampling (SIR), i.e., particle filtering. Specifically, the belief $\mathcal{B}_{\tau}$ is represented by a set of weighted particles $\mathcal{B}_{\tau}=\{w^{(i)}_{\tau},X^{(i)}_{\tau}\}_{i=1}^{N_{s}}$, where $X^{(i)}_{\tau}=[x^{1}_{\tau},\ldots,x^{N}_{\tau}]^{\top}$. These particles are propagated to the next time step according to the process dynamics and reweighted using the likelihood function to compute the posterior. For notational convenience, we will often write $t+\tau|t$ as $\tau$ when no ambiguity arises. The functional $\mathcal{C}_{t+\tau|t}:\mathcal{B}_{t+\tau|t}(X_{t+\tau|t}|Z_{1:t+\tau|t})\rightarrow[0,1]$ in Eq. (6a) quantifies the uncertainty of the fire-front state $X_{t+\tau|t}$ encoded in the posterior distribution at time $t+\tau|t$ , conditioned on all hypothetical measurements $Z_{1:t+\tau|t}$. This uncertainty is measured by the Risk-Weighted Dispersion (RWD) defined as:
$\mathcal{C}_{t+\tau|t}(\mathcal{B}_{t+\tau|t})=\frac{1}{\omega}\sum_{\varepsilon\in\tilde{\mathcal{E}}}\mathcal{R}(\varepsilon)\det(\Sigma^{\varepsilon}_{t+\tau|t}),$ (7)
@@
where the environment $\mathcal{E}$ is discretized in space to form a 2D grid $\tilde{\mathcal{E}}$, composed of a finite number of non-overlapping, equally sized cells $\tilde{\mathcal{E}}=\{\varepsilon_{1},\ldots,\varepsilon_{|\tilde{\mathcal{E}}|}\}$, such that $\bigcup_{i=1}^{|\tilde{\mathcal{E}}|}\varepsilon_{i}=\mathcal{E}$. The term $\mathcal{R}(\varepsilon)\in[0,1]$ denotes the risk value associated with cell $\varepsilon$, reflecting the severity of fire presence in that region. The random quantity $\det(\Sigma_{t+\tau|t}^{\varepsilon})$ is the determinant of the sample covariance matrix $\Sigma_{t+\tau|t}^{\varepsilon}$ computed from all particle points residing in cell $\varepsilon$ at time step $t+\tau|t$, and $\omega$ is a scaling factor ensuring that $\mathcal{C}_{t+\tau|t}(\mathcal{B}_{t+\tau|t})\in[0,1]$. Finally, the parameter $\nu\in(0,1]$ in Eq. (6a) is a discount factor that controls the relative importance of future decisions.
@@ III-C Adaptive LCB-guided Policy Search
where $s^{\pi}_{t+\tau|t}$ is the state encountered at level $\tau$ in the horizon under policy $\pi$ when starting from state $s_{t}$. Subsequently, the optimal policy that minimizes the expected cumulative cost over the horizon is given by $\pi^{\star}=\arg\min_{\pi\in\Pi_{T}}\,V^{\pi}_{t}(s_{t})$. To solve this problem we build upon the Upper Confidence Bound 1 (UCB1) framework [2] and we treat each sequence $\pi=\{\hat{u}_{t+\tau-1|t}\}_{\tau=1}^{T}\in\mathbb{U}^{T}$ as an arm in a multi-armed bandit. We use rollouts to simulate policy outcomes and employ a UCB1-like adaptive selection strategy to efficiently balance exploration and exploitation.
@@
4:  Sample policy: $\hat{\pi}\!=\!\arg\min\mathrm{LCB}^{n}(\pi,s_{t})$ (Eq. (9))
5:  for $\tau=1,\dots,T$ do
6:   Compute predictive density: $\mathcal{B}^{-}_{t+\tau|t}$ via Eq. (6b)
7:   Move to new state: $\hat{y}_{t+\tau|t}=f_{a}\left(\hat{y}_{t+\tau-1|t},\hat{\pi}_{\tau}\right)$
8:   Sample meas.: $Z_{t+\tau|t}\sim p(Z|X_{t+\tau|t},\hat{y}_{t+\tau|t})$
9:   Compute posterior: $\mathcal{B}_{t+\tau|t}$ via Eq. (6f)
10:   Compute stage cost: $c_{\tau}=c_{\tau-1}+\nu^{\tau}\mathcal{C}_{\tau}(\mathcal{B}_{t+\tau|t})$
11:  end for
12:  Update LCB: $Q^{n}(\hat{\pi},s_{t})=\mathrm{mean}\left(Q^{n-1}(\hat{\pi},s_{t}),c_{T}\right)$
13:                           $I^{n}(\hat{\pi})=I^{n}(\hat{\pi})+1$
14:end for
15:Output: Optimal policy $\pi^{\star}=\arg\min~\mathrm{LCB}^{n_{\text{max}}}(\pi,s_{t})$
@@
The proposed adaptive-search algorithm iteratively computes a Lower Confidence Bound (LCB) on the expected total cost associated with each $T$-finite control sequence, which is then used to adaptively select the next policy to simulate. Specifically, for an agent at state $s_{t}$, we define the LCB score of a policy $\pi\in\Pi_{T}$ at iteration $n$ as:
$\mathrm{LCB}^{n}(\pi,s_{t})=\begin{cases}Q^{n}(\pi,s_{t})-\sqrt{\frac{2\ln n}{I^{n}(\pi)}},&\text{if }I^{n}(\pi)\neq 0,\\ \mathrm{LCB}^{n}_{\min},&\text{otherwise},\end{cases}$ (9)
@@
where $I^{n}(\pi)$ denotes the number of times policy $\pi$ has been simulated up to iteration $n$, and $Q^{n}(\pi,s_{t})=\frac{1}{I^{n}(\pi)}\sum_{i=1}^{n}\delta(\pi^{i}-\pi)L(\pi,s_{t})$ is the sample mean of the cumulative cost incurred by policy $\pi$. Here, $L(\pi,s_{t})=\sum_{\tau=1}^{T}\nu^{\tau}\mathcal{C}_{\tau}\left(s^{\pi}_{t+\tau|t}\right)$ is the total discounted cost of executing policy $\pi$ starting from state $s_{t}$, and $\delta(\cdot)$ is the discrete Dirac delta function. The constant $\mathrm{LCB}^{n}_{\min}<-\sqrt{2\ln n}$ ensures that any policy not yet selected by iteration $n$ will have an LCB value lower than that of any previously selected policy, thereby guaranteeing it will be explored with higher probability in subsequent iterations. Finally, the term $\sqrt{\frac{2\ln n}{I^{n}(\pi)}}$ serves as an exploration bonus, encouraging the selection of under-explored policies. Under the standard UCB1 assumptions, the expected fraction of iterations on which LCB selects a suboptimal policy is $O\!\big(\tfrac{\ln n}{n}\big)$; hence the optimal policy is selected with asymptotic frequency $1$ in expectation as $n\to\infty$.
@@
The complete LCB-guided adaptive search algorithm is shown in Algorithm 1. At each time step $t$, the algorithm identifies the optimal policy $\pi^{\star}$ over the horizon $\{t+\tau|t\}_{\tau=1}^{T}$. The agent then executes the first control input of $\pi^{\star}$, transitions to a new state, receives the “real” measurement, computes the posterior belief $\mathcal{B}_{t+1}$, and the algorithm is re-applied at the next time step $t+1$. We should note that intelligent pruning techniques, such as $\epsilon$-suboptimal reductions [1], can be designed and integrated into the proposed approach to focus on the most promising set of control inputs at each time step thereby reducing the runtime complexity.

### v2

@@ III-B Information-seeking Predictive Control
For each prediction time step $\tau\in\{1,\ldots,T\}$ in the horizon, the agent predicts the fire front state $X_{\tau}$ forward in time using the Bayesian prediction step, based on the transition density $p(X_{\tau}|X_{\tau-1})$ and the (pseudo) posterior belief from the previous time step, $\mathcal{B}_{t+\tau-1|t}(X_{\tau-1}|Z_{1:t+\tau-1|t})$, as shown in Eqs. (13b)-(13c). The constraints in Eqs. (13d)-(13e) arise from the agent dynamics, which are assumed to be deterministic in this work. At the predicted time step $\tau$, given the agent’s predicted state $y_{t+\tau|t}$ and sensing range $O_{t+\tau|t}$, the agent receives the predicted measurement set $Z_{t+\tau|t}$. This set is then used to compute the posterior belief $\mathcal{B}_{t+\tau|t}(X_{\tau}|Z_{1:t+\tau|t})$ via the Bayesian correction step, as shown in Eq. (13f), using the joint likelihood function $p(Z_{t+\tau|t}|X_{\tau},y_{t+\tau|t})$ and the predicted density. This posterior distribution subsequently becomes the prior for the next prediction step, continuing the recursive process. The predicted measurements $Z_{t+\tau|t}$ represent hypothetical observations based on the predicted control inputs and the anticipated fire front state. Since actual measurements are only available after executing the control actions, the objective in Eq. (13a) requires taking an expectation over all possible future measurement sequences. This enables informed decision-making by accounting for potential outcomes without executing the corresponding trajectories.
@@
Equations (13b)-(13f) admit no closed form: the model is nonlinear and non-Gaussian with set-valued (multi-object, multi-measurement) observations, so Kalman-type filters are inapplicable. Therefore, the recursion is implemented using Sequential Importance Resampling (SIR), i.e., particle filtering. Specifically, the belief $\mathcal{B}_{\tau}$ is represented by a set of weighted particles $\mathcal{B}_{\tau}=\{w^{(i)}_{\tau},X^{(i)}_{\tau}\}_{i=1}^{N_{s}}$, where $X^{(i)}_{\tau}=[x^{1}_{\tau},\ldots,x^{N}_{\tau}]^{\top}$. These particles are propagated to the next time step according to the process dynamics and reweighted using the likelihood function to compute the posterior. For notational convenience, we will often write $t+\tau|t$ as $\tau$ when no ambiguity arises. The functional $\mathcal{C}_{t+\tau|t}:\mathcal{B}_{t+\tau|t}(X_{t+\tau|t}|Z_{1:t+\tau|t})\rightarrow[0,1]$ in Eq. (13a) quantifies the uncertainty of the fire-front state $X_{t+\tau|t}$ encoded in the posterior distribution at time $t+\tau|t$, conditioned on all hypothetical measurements $Z_{1:t+\tau|t}$. This uncertainty is measured by the Risk-Weighted Dispersion (RWD), defined as:
$\mathcal{C}_{t+\tau|t}\!\big(\mathcal{B}_{t+\tau|t}\big)=\frac{1}{\omega}\sum_{\varepsilon\in\tilde{\mathcal{E}}}\mathcal{R}(\varepsilon)\,\det\!\big(\Sigma^{\varepsilon}_{t+\tau|t}\big),$ (14)
@@
where the environment $\mathcal{E}$ is discretized in space to form a 2D grid $\tilde{\mathcal{E}}$, composed of a finite number of non-overlapping, equally sized cells $\tilde{\mathcal{E}}=\{\varepsilon_{1},\ldots,\varepsilon_{|\tilde{\mathcal{E}}|}\}$, such that $\bigcup_{i=1}^{|\tilde{\mathcal{E}}|}\varepsilon_{i}=\mathcal{E}$. The term $\mathcal{R}(\varepsilon)\in[0,1]$ denotes the risk value associated with cell $\varepsilon$, reflecting the severity of fire presence in that region. The random quantity $\det(\Sigma_{t+\tau|t}^{\varepsilon})$ is the determinant of the sample covariance matrix $\Sigma_{t+\tau|t}^{\varepsilon}$ computed from all particle points residing in cell $\varepsilon$ at time step $t+\tau|t$, and $\omega$ is a scaling factor ensuring that $\mathcal{C}_{t+\tau|t}\!\big(\mathcal{B}_{t+\tau|t}\big)\in[0,1]$.
In practice, the computation of the aggregate cell dispersion $\det\!\big(\Sigma^{\varepsilon}_{t+\tau|t}\big)$ is implemented via a vertex-centric pooling operation. For each fire-front vertex $i\in\{1,\dots,N\}$, we explicitly compute its $2\times 2$ covariance matrix $\Sigma^{i}_{t+\tau|t}$ across its entire ensemble of $N_{s}$ particle realizations, which captures the full spatial dispersion of the vertex. Subsequently, the cell dispersion term $\det\!\big(\Sigma^{\varepsilon}_{t+\tau|t}\big)$ in Eq. (14) is evaluated operationally by summing the individual uncertainty volumes, i.e., $\sum_{i}\det\!\big(\Sigma^{i}_{t+\tau|t}\big),$ over all vertices whose expected spatial location (MMSE) resides within cell $\varepsilon$.
Finally, the parameter $\nu\in(0,1]$ in Eq. (13a) is a discount factor that controls the relative importance of future decisions.
@@ III-C Adaptive LCB-guided Policy Search
where $s^{\pi}_{t+\tau|t}$ is the state encountered at level $\tau$ in the horizon under policy $\pi$ when starting from state $s_{t}$. Subsequently, the optimal policy that minimizes the expected cumulative cost over the horizon is given by
$\pi^{\star}=\arg\min_{\pi\in\Pi_{T}}\,V^{\pi}_{t}(s_{t})$ (16)
Observe that the optimization problem in Eq. (16) is equivalent to Problem (P1) under the assumption of a finite number of admissible control inputs. Solving this equivalent MDP formulation however, introduces its own set of challenges. In particular, the continuous belief-space represented by particles and the stochastic nature of the measurements result in an effectively infinite state space. This makes classical dynamic programming (DP) methods, such as value iteration and backward induction, infeasible. These approaches rely on a finite state space to compute value functions backward from the planning horizon and therefore do not apply here. Furthermore, the transition dynamics are not explicitly known in closed form, nor can they be represented in a tabular format.
To solve this problem we build upon the Upper Confidence Bound 1 (UCB1) framework [2] and we treat each sequence $\pi=\{\hat{u}_{t+\tau-1|t}\}_{\tau=1}^{T}\in\mathbb{U}^{T}$ as an arm in a multi-armed bandit. We use rollouts to simulate policy outcomes and employ a UCB1-like adaptive selection strategy to efficiently balance exploration and exploitation. In particular, the UCB1 algorithm iteratively computes an upper confidence bound score on the expected reward for each arm (where arms are considered the actions), by adding the sample mean reward of the arm to an exploration bonus that depends on both the total number of arm pulls and the number of pulls for that specific arm.
@@
4:  Sample policy: $\hat{\pi}\!=\!\arg\min\mathrm{LCB}^{n-1}(\pi,s_{t})$, Eq. (17)
5:  Initialize rollout cost: $c_{0}=0$
6:  for $\tau=1,\dots,T$ do
7:   Compute predictive density: $\mathcal{B}^{-}_{t+\tau|t}$ via Eq. (13b)
8:   Move to new state: $\hat{y}_{t+\tau|t}=f_{a}\left(\hat{y}_{t+\tau-1|t},\hat{\pi}_{\tau}\right)$
9:   Sample latent state: $\tilde{X}_{t+\tau|t}\sim\mathcal{B}^{-}_{t+\tau|t}$
10:   Sample meas.: $Z_{t+\tau|t}\sim p(Z|\tilde{X}_{t+\tau|t},\hat{y}_{t+\tau|t})$
11:   Compute posterior: $\mathcal{B}_{t+\tau|t}$ via Eq. (13f)
12:   Compute stage cost: $c_{\tau}=c_{\tau-1}+\nu^{\tau}\mathcal{C}_{\tau}(\mathcal{B}_{t+\tau|t})$
13:  end for
14:  Update: $Q^{n}(\hat{\pi},s_{t})\!=\!\mathrm{mean}\left(Q^{n-1}(\hat{\pi},s_{t}),\frac{c_{T}}{L_{\max}}\right)$
15:                  $I^{n}(\hat{\pi})=I^{n-1}(\hat{\pi})+1$
16:end for
17:Output: Optimal policy $\pi^{\star}=\arg\min~Q^{n_{\text{max}}}(\pi,s_{t})$
This formulation is grounded in the “optimism under uncertainty” principle and leverages concentration inequalities (i.e., Chernoff-Hoeffding bounds) to guarantee near-optimal regret (i.e., the expected loss between the optimal policy and selected policies). Based on these scores, the policy always selects the arm with the highest current upper confidence bound, which enables asymptotic convergence to the optimal policy [2].
@@
Specifically, the proposed adaptive-search algorithm iteratively computes a Lower Confidence Bound (LCB) on the expected total cost associated with each $T$-finite control sequence, which is then used to adaptively select the next policy to simulate. Specifically, for an agent at state $s_{t}$, we define the LCB score of a policy $\pi\in\Pi_{T}$ at iteration $n$ as:
$\mathrm{LCB}^{n-1}(\pi,s_{t})\!\!=\!\!\begin{cases}\!\!Q^{n-1}(\pi,s_{t})\!-\!\sqrt{\frac{2\ln n}{I^{n-1}(\pi)}},\!\!\!&I^{n-1}(\pi)>0,\\ \!\!\mathrm{LCB}^{n-1}_{\min},\!\!\!&\text{otherwise},\end{cases}$ (17)
@@
where $I^{n-1}(\pi)$ denotes the number of times policy $\pi$ has been simulated prior to iteration $n$, and $Q^{n-1}(\pi,s_{t})=\frac{1}{I^{n-1}(\pi)}\sum_{i=1}^{n-1}\mathds{1}_{\{\hat{\pi}^{i}=\pi\}}L^{(i)}(\pi,s_{t})$ is the empirical sample mean of the cumulative cost incurred by policy $\pi$ up to iteration n1n-1. Here, $\hat{\pi}^{i}$ is the policy selected at iteration $i$, $\mathds{1}_{\{\cdot\}}$ is the indicator function, and $L^{(i)}(\pi,s_{t})=\frac{1}{L_{\max}}\sum_{\tau=1}^{T}\nu^{\tau}\mathcal{C}_{\tau}\!\left(s^{\pi,(i)}_{t+\tau|t}\right)$ is the normalized total discounted cost realized explicitly during the $i$-th simulated rollout of policy $\pi$ starting from state $s_{t}$, with $L_{\max}=\sum_{\tau=1}^{T}\nu^{\tau}$. The constant $\mathrm{LCB}^{n-1}_{\min}<-\sqrt{2\ln n}$ ensures that any policy not yet selected by iteration $n$ will have an LCB value lower than that of any previously selected policy, thereby guaranteeing it will be explored with higher probability in subsequent iterations. Finally, the term $\sqrt{\frac{2\ln n}{I^{n-1}(\pi)}}$ serves as an exploration bonus, encouraging the selection of under-explored policies.
@@
The complete LCB-guided adaptive search algorithm is shown in Algorithm 1. At each time step $t$, the algorithm identifies the optimal policy $\pi^{\star}$ over the horizon $\{t+\tau|t\}_{\tau=1}^{T}$. The agent then executes the first control input of $\pi^{\star}$, transitions to a new state, receives the “real” measurement, computes the posterior belief $\mathcal{B}_{t+1}$, and the algorithm is re-applied at the next time step $t+1$. In Note that in Alg. 1, $\mathrm{mean}\!\big(Q^{n-1}(\hat{\pi},s_{t}),\frac{c_{T}}{L_{\max}}\big)$ stands for the recursive empirical average update, computed as $Q^{n-1}(\hat{\pi},s_{t})+\frac{1}{I^{n}(\hat{\pi})}\!\left(\frac{c_{T}}{L_{\max}}-Q^{n-1}(\hat{\pi},s_{t})\right)$. Finally, in Alg. 1, the predicted measurements are simulated via ancestral sampling i.e., a latent fire-front state is first drawn from the predictive belief $\tilde{X}_{t+\tau|t}\sim\mathcal{B}^{-}_{t+\tau|t}$, and the measurement set is subsequently generated from the measurement likelihood conditioned on this latent state and the planned agent state. We should note that intelligent pruning techniques, such as $\epsilon$-suboptimal reductions [1], can be designed and integrated into the proposed approach to focus on the most promising set of control inputs at each time step thereby reducing the runtime complexity.
Theorem 1 (Completeness): Let $\Pi_{T}$ be the set of all candidate policies in the LCB framework. Then for every optimal policy $\pi^{\star}\in\Pi_{T}$, there exists iteration $n$ such that $\pi^{\star}$ is selected at least once by the algorithm.
Proof:
By construction of the LCB algorithm, each policy $\pi\in\Pi_{T}$ is initialized in a manner that guarantees it is tried at least once. Thus, any policy that has never been tried up to a certain point will have a lower LCB score than those policies with a finite sample count, ensuring it is selected at least once. Hence $\pi^{\star}$ is included among the solutions eventually tried. ∎
Theorem 2 (Asymptotic Convergence): Suppose there is a unique optimal policy $\pi^{\star}$ whose expected cost is strictly smaller than that of any other $\pi\in\Pi_{T}$. Under LCB, the expected average regret decays on the order of $O\bigl(\tfrac{\ln n}{n}\bigr)$, and the fraction of times $\pi^{\star}$ is selected converges to $1$ as $n\to\infty$.
Proof:
By Theorem 1, every policy in $\Pi_{T}$ is selected at least once under LCB. Moreover, LCB follows the same selection rule as UCB1 for minimization problems, adapted to entire control sequences, effectively treating each policy as a bandit arm with a stationary cost distribution and stage costs bounded in the range $[0,1]$. Consequently, the classical finite-time analysis on UCB1 [2] applies directly: the expected number of times any suboptimal policy is selected is bounded by $O(\ln n)$. Hence, the expected fraction of times a suboptimal policy is chosen decreases as $O\bigl(\frac{\ln n}{n}\bigr)$, which implies the expected average regret shrinks at that same rate. Thus, as $n\to\infty$, the average regret approaches 0 and the fraction of times the optimal policy $\pi^{\star}$ is selected asymptotically converges to 1. ∎